PERMUTATIONS OF SET
Calculate each of the following permutations.
1. 6P2______ 2. 8P4______ 3. 9P1_______ 4.7P4_______ 5. 8P6______ 6. 10P7______
7. 15P3________ 8. 13P9_______ 9. 20P5_________


Answer by Edwin McCravy(20086)

PERMUTATIONS OF SET
Calculate each of the following permutations.

1. 6P2______ 

Start with 6, subtract 1 each time until you have 2 factors,
then multiply them all together.

6*5 = 30 

There are 30 different ways to pick 2 things from 6 and arrange
them in a row.

------------------------

2. 8P4______ 

Start with 8, subtract 1 each time until you have 4 factors,
then multiply them all together.

8*7*6*5 = 1680

There are 1680 different ways to pick 4 things from 8 and arrange
them in a row.

------------------------

3. 9P1_______ 

Start with 9.  You can't subtract 1 for you already have 1 factor 
from the start.  So the answer is just the 1 factor 9

There are 9 different ways to pick 1 thing from 9.

------------------------


4. 7P4_______ 

Start with 7, subtract 1 each time until you have 4 factors,
then multiply them all together.

7*6*5*4 = 840

There are 840 different ways to pick 4 things from 7 and arrange
them in a row.

------------------------

5. 8P6______ 

Start with 8, subtract 1 each time until you have 6 factors,
then multiply them all together.

8*7*6*5*4*3 = 20160

There are 20160 different ways to pick 6 things from 8 and arrange
them in a row.

------------------------

6. 10P7______

Start with 10, subtract 1 each time until you have 7 factors,
then multiply them all together.

10*9*8*7*6*5*4 = 604800

There are 604800 different ways to pick 7 things from 10 and arrange
them in a row.

------------------------

7. 15P3________ 

Start with 15, subtract 1 each time until you have 3 factors,
then multiply them all together.

15*14*13 = 2730

There are 2730 different ways to pick 3 things from 15 and arrange
them in a row.

------------------------

8. 13P9_______ 

Start with 13, subtract 1 each time until you have 9 factors,
then multiply them all together.

13*12*11*10*9*8*7*6*5 = 259459200

There are 259459200 different ways to pick 9 things from 13 and arrange
them in a row.

------------------------

9. 20P5_________

Start with 20, subtract 1 each time until you have 5 factors,
then multiply them all together.

20*19*18*17*16 = 1860480

There are 1860480 different ways to pick 5 things from 20 and arrange
them in a row.

------------------------

Edwin